### v1

# Enhanced Muonization by Active-Sterile Neutrino Mixing in Protoneutron Stars

Abstract
We study $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the protoneutron star (PNS) created in a core-collapse supernova (CCSN). We point out the importance of the feedback on the general composition of the PNS in addition to the obvious feedback on the $\nu_{\mu}$ lepton number. We show that for our adopted mixing parameters $\delta m^{2}\sim 10^{2}$ keV2 and $\sin^{2}2\theta$ consistent with the current constraints, sterile neutrino production is dominated by the Mikheyev–Smirnov–Wolfenstein conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ and that the subsequent escape of $\bar{\nu}_{s}$ increases the $\nu_{\mu}$ lepton number, which in turn enhances muonization of the PNS primarily through $\nu_{\mu}+n\to p+\mu^{-}$. We find that this enhancement is amplified by diffusion of the $\nu_{e}$ and $\nu_{\mu}$ lepton numbers. While these results are qualitatively robust, their quantitative effects on the dynamics and active neutrino emission of CCSNe should be evaluated by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the simulations.
preprint: N3AS-24-013
I Introduction
Sterile neutrinos ($\nu_{s}$) associated with a vacuum mass eigenstate of keV-scale mass are a viable candidate for the mysterious dark matter Kusenko:2009up ; Drewes:2016upu ; Abazajian:2017tcc ; Boyarsky:2018tvu ; Dasgupta:2021ies . Because of their tiny mixing with active neutrinos ($\nu_{\alpha}$), they can decay radiatively, $\nu_{s}\to\nu_{\alpha}+\gamma$, thereby allowing experimental probes through the decay X-ray line. In fact, major experimental constraints on active-sterile neutrino mixing primarily come from such X-ray searches Abazajian:2001vt ; Boyarsky:2007ge ; Ng:2019gch ; Roach:2019ctw ; Foster:2021ngm ; Roach:2022lgo . There are also other significant constraints from phase-space considerations (the Tremaine-Gunn bound) PhysRevLett.42.407 ; Boyarsky:2008ju ; Gorbunov:2008ka ; DiPaolo:2017geq , measurement of Lyman-$\alpha$ forests Boyarsky:2008xj ; Irsic:2017ixq ; Zelko:2022tgf , and terrestrial nuclear decay searches Smith:2016vku ; KATRIN:2018oow ; Benso:2019jog ; Balantekin:2023jlg . Among many theoretical explorations, a number of studies have investigated the effects of active-sterile neutrino mixing in core-collapse supernovae (CCSNe) Shi:1993ee ; Nunokawa:1997ct ; Hidaka:2006sg ; Hidaka:2007se ; 1993APh…..1..165R ; 2011PhRvD..83i3014R ; Warren:2014qza ; Warren:2016slz ; 2019PhRvD..99d3012A ; 2019JCAP…12..019S ; Suliga:2020vpz ; PhysRevD.106.015017 ; PhysRevD.108.063025 ; Carenza:2023old ; Chauhan:2023sci . In this paper, we focus on the mixing between $\nu_{\mu}$ ($\bar{\nu}_{\mu}$) and $\nu_{s}$ ($\bar{\nu}_{s}$) with a vacuum mass-squared difference of $\delta m^{2}\sim 10^{2}$ keV2 in protoneutron stars (PNSs) created in CCSNe. In particular, we treat this mixing in the presence of muons. In contrast to Ref. PhysRevD.106.015017 , which focused on the energy loss due to $\nu_{x}$-$\nu_{s}$ and $\bar{\nu}_{x}$-$\bar{\nu}_{s}$ ($x=\mu$, $\tau$) mixing and stated that the presence of muons makes little difference, we show that inclusion of muons leads to more complicated feedback effects for $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing than those for $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing.
Due to the high densities and temperatures in the PNS, muon production or muonization (e.g., $e^{-}+\bar{\nu}_{e}\to\mu^{-}+\bar{\nu}_{\mu}$, $\nu_{\mu}+n\to p+\mu^{-}$) is energetically allowed and dynamically important for the neutrino-driven explosion 2017PhRvL.119x2702B . The dynamic effects of muonization can be understood qualitatively as follows. For the conditions in the PNS, electrons are highly relativistic, and their contribution to the pressure is $\approx 1/3$ of their energy density. Without muons, charge neutrality requires an equal number of electrons and protons. For a given initial distribution of protons, as some electrons are converted into muons, a fraction of the electron energy is converted into the rest mass energy of muons, thereby reducing the net pressure. Consequently, the PNS contracts faster and the resulting increase in neutrino luminosity enhances heating of the material outside the PNS, which helps the neutrino-driven explosion 2017PhRvL.119x2702B . The conditions in the PNS also facilitate conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ through the Mikheyev–Smirnov–Wolfenstein (MSW) effect MikheyeSmirnov1985 ; 1978PhRvD..17.2369W , which tends to increase the $\nu_{\mu}$ lepton number, thereby enhancing muonization mainly through the reaction $\nu_{\mu}+n\to p+\mu^{-}$. While the corresponding dynamic effects must be investigated by incorporating $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in CCSN simulations, our goal in this paper is to illustrate how such mixing enhances muonization in the PNS and to estimate the potential enhancement.
The rest of the paper is organized as follows. In Section II we discuss the production of sterile neutrinos through $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. Much of the treatment is similar to that of $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing, and we follow the discussion in Ref. PhysRevD.108.063025 . In Section III we discuss the treatment of the feedback effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the presence of muons. In Section IV we present example calculations to illustrate the effects of such mixing on muonization in the PNS. In Section V we discuss our results and give conclusions.
II Production of sterile neutrinos in a PNS
We assume spherical symmetry, for which the conditions in the PNS are functions of radius $r$ and time $t$ only. We focus on the region where all active neutrinos are diffusing and assume that all particles of concern, $n$, $p$, $e^{\pm}$, $\mu^{\pm}$, $\nu_{e}$, $\bar{\nu}_{e}$, $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{\tau}$, and $\bar{\nu}_{\tau}$, have Fermi-Dirac energy distributions characterized by the same temperature $T$ but species-specific chemical potentials. For example, the $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ energy distributions (number densities per unit energy interval per unit solid angle) are
$\frac{d^{2}n_{\nu_{\mu}}}{dEd\Omega}$ $=\frac{1}{(2\pi)^{3}}\frac{E^{2}}{\exp[(E-\mu_{\nu_{\mu}})/T]+1},$ (1a)
$\frac{d^{2}n_{\bar{\nu}_{\mu}}}{dEd\Omega}$ $=\frac{1}{(2\pi)^{3}}\frac{E^{2}}{\exp[(E+\mu_{\nu_{\mu}})/T]+1},$ (1b)
where $\mu_{\nu_{\mu}}$ is the $\nu_{\mu}$ chemical potential and we have used $\mu_{\bar{\nu}_{\mu}}=-\mu_{\nu_{\mu}}$. The corresponding $\nu_{\mu}$ lepton number fraction (net number per baryon) is
$Y_{\nu_{\mu}}=\frac{n_{\nu_{\mu}}-n_{\bar{\nu}_{\mu}}}{n_{b}}=\frac{T^{3}\eta_{\nu_{\mu}}}{6n_{b}}\left(1+\frac{\eta_{\nu_{\mu}}^{2}}{\pi^{2}}\right),$ (2)
where $\eta_{\nu_{\mu}}=\mu_{\nu_{\mu}}/T$ and $n_{b}$ is the baryon number density.
The treatment of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing would mirror that of $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing (e.g., 1993APh…..1..165R ; 2011PhRvD..83i3014R ; 2019PhRvD..99d3012A ; 2019JCAP…12..019S ; PhysRevD.106.015017 ; PhysRevD.108.063025 ) were there no muons. With muons, the effective potential for $\nu_{\mu}$-$\nu_{s}$ mixing is
$V_{\nu}=\sqrt{2}G_{F}n_{b}\left(-\frac{Y_{n}}{2}+Y_{\mu}+Y_{\nu_{e}}+2Y_{\nu_{\mu}}\right),$ (3)
where $G_{F}$ is the Fermi constant, and $Y_{\alpha}$ ($\alpha=n$, $\mu$, $\nu_{e}$, and $\nu_{\mu}$) is the net number fraction for species $\alpha$. In Eq. (3), we have applied the constraint of charge neutrality, $Y_{e}+Y_{\mu}=Y_{p}$, so the contributions from forward scattering of $\nu_{\mu}$ on protons and electrons do not appear explicitly. We have also assumed that there is no net $\nu_{\tau}$ lepton number in the PNS (i.e., $Y_{\nu_{\tau}}=0$ and $\mu_{\nu_{\tau}}=0$). Note that $V_{\nu}$, $n_{b}$, and $Y_{\alpha}$ are all functions of $r$ and $t$. For convenience, here and below, we usually suppress such radial and temporal dependence. The potential for $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing is $V_{\bar{\nu}}=-V_{\nu}$.
The potential $V_{\nu}$ ($V_{\bar{\nu}}$) modifies the mixing angle between $\nu_{\mu}$ ($\bar{\nu}_{\mu}$) and $\nu_{s}$ ($\bar{\nu}_{s}$) in the PNS. For $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ with energy $E$, the effective mixing angles are given by
$\sin^{2}2\theta_{\nu}$ $=\frac{\Delta^{2}\sin^{2}2\theta}{(\Delta\cos 2\theta-V_{\nu})^{2}+\Delta^{2}\sin^{2}2\theta},$ (4a)
$\sin^{2}2\theta_{\bar{\nu}}$ $=\frac{\Delta^{2}\sin^{2}2\theta}{(\Delta\cos 2\theta-V_{\bar{\nu}})^{2}+\Delta^{2}\sin^{2}2\theta},\,$ (4b)
where $\theta\ll 1$ is the vacuum mixing angle and $\Delta=\delta m^{2}/(2E)$. For the conditions in the PNS, $V_{\nu}<0$ and $V_{\bar{\nu}}>0$. Therefore, $\nu_{\mu}$-$\nu_{s}$ mixing is suppressed whereas $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing can be enhanced by the MSW resonance. This resonance occurs when $\Delta\cos 2\theta=V_{\bar{\nu}}$, and defines a resonance energy $E_{R}=\delta m^{2}\cos 2\theta/(2V_{\bar{\nu}})$.
Resonant conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ is the dominant mechanism for producing sterile neutrinos inside the PNS (e.g., 2019PhRvD..99d3012A ; 2019JCAP…12..019S ; PhysRevD.106.015017 ; PhysRevD.108.063025 ). We follow Ref. PhysRevD.108.063025 and adopt the following rate of change in $Y_{\nu_{\mu}}$ due to the MSW conversion of $\bar{\nu}_{\mu}$ and the subsequent escape of $\bar{\nu}_{s}$ from the PNS:
$\dot{Y}_{\nu_{\mu}}^{\rm MSW}=\Theta(\lambda_{R}-\delta r)\frac{\pi E_{R}(1-P_{\rm LZ}^{2})}{n_{b}H_{R}}\left.\frac{d^{2}n_{\bar{\nu}_{\mu}}}{dEd\Omega}\right|_{E_{R}}.$ (5)
In the above equation, $\Theta(x)$ is the Heaviside step function, $\lambda_{R}=1/[n_{b}\sigma(E_{R})]$ is the mean free path at the resonant neutrino energy $E_{R}$, $\sigma(E)\approx G_{F}^{2}E^{2}/\pi$ is the cross section governing $\bar{\nu}_{\mu}$ diffusion (e.g., 2011PhRvD..83i3014R ), $\delta r=2H_{R}\tan 2\theta$ is the width of the resonance region, $H_{R}$ is the scale height $|\partial\ln V_{\bar{\nu}}/\partial r|_{E_{R}}^{-1}$ with the derivative taken at the resonance radius for $E_{R}$, and
$P_{\rm LZ}=\exp\left(-\frac{\pi\delta m^{2}H_{R}\sin^{2}2\theta}{4E_{R}\cos 2\theta}\right)\,,$ (6)
is the Landau-Zener survival probability for a radially-propagating $\bar{\nu}_{\mu}$ after it crosses the resonance. Note that a radially-outgoing $\bar{\nu}_{\mu}$ crosses the resonance only once, whereas a radially-incoming $\bar{\nu}_{\mu}$ crosses the resonance twice (see Fig. 1). Both situations are taken into account PhysRevD.108.063025 by Eq. (5).
Figure 1: Illustration of two types of escaping $\bar{\nu}_{s}$. One type is converted from radially outgoing $\bar{\nu}_{\mu}$ that experience a single MSW resonance, while the other type is converted from radially incoming $\bar{\nu}_{\mu}$ that experience two resonances (with conversion occurring only at the first resonance).
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Comparison of the radial profiles of $T$, $\rho$, $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ at $t_{\rm pb}=1$ and 1.25 s for a $20\,M_{\odot}$ CCSN model (with the SFHo nuclear equation of state and with muonization) provided by the Garching group 2017PhRvL.119x2702B ; Bollig2016 .
Apart from the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$, $\nu_{s}$ ($\bar{\nu}_{s}$) can be produced via collisions of $\nu_{\mu}$ ($\bar{\nu}_{\mu}$) with the PNS constituents. In this scenario, $\nu_{\mu}$ ($\bar{\nu}_{\mu}$) evolves as a linear combination of two effective mass eigenstates between collisions. Upon collision, which is predominantly with baryons, the wave function collapses, and a $\nu_{s}$ ($\bar{\nu}_{s}$) is produced with a probability proportional to $\sin^{2}2\theta_{\nu}$ $(\sin^{2}2\theta_{\bar{\nu}})$. For the mixing parameters of interest, the collisional production of $\nu_{s}$ and $\bar{\nu}_{s}$ inside a PNS is much less efficient than the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. Nevertheless, for completeness, we also include the following contribution to the rate of change in $Y_{\nu_{\mu}}$ from the collisional production of $\nu_{s}$ and $\bar{\nu}_{s}$:
$\dot{Y}_{\nu_{\mu}}^{\rm coll}$ $=G_{F}^{2}\left(\int_{E_{\bar{\nu}_{\mu}}}dE\,E^{2}\sin^{2}2\theta_{\bar{\nu}}\frac{d^{2}n_{\bar{\nu}_{\mu}}}{dEd\Omega}\right.$
$\left.-\int_{0}^{\infty}dE\,E^{2}\sin^{2}2\theta_{\nu}\frac{d^{2}n_{\nu_{\mu}}}{dEd\Omega}\right)\,.$ (7)
In the above equation, the integration over $E_{\bar{\nu}_{\mu}}$ excludes the range $[E_{R}-\delta E/2,E_{R}+\delta E/2]$ with $\delta E=(\lambda_{R}/H_{R})E_{R}$ when $\lambda_{R}$ exceeds $\delta r$ and MSW conversion occurs for resonant $\bar{\nu}_{\mu}$, but covers the entire energy range otherwise.
III Feedback from active-sterile neutrino mixing
We focus on the region where active neutrinos are trapped. Because sterile neutrinos can escape from this region, $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing leads to changes of the $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ energy distributions, which in turn cause evolution of $Y_{\nu_{\mu}}$, and hence that of $V_{\nu}$ and $V_{\bar{\nu}}$. This type of feedback also occurs for $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing 2011PhRvD..83i3014R ; 2019JCAP…12..019S ; PhysRevD.106.015017 ; PhysRevD.108.063025 . What is new here, however, is that the changes of the $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ energy distributions cause evolution of $Y_{n}$, $Y_{p}$, and $Y_{\mu}$ through the reactions
$\nu_{\mu}+n$ $\to p+\mu^{-},$ (8a)
$\bar{\nu}_{\mu}+p$ $\to n+\mu^{+}.$ (8b)
Therefore, the feedback effects for $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing are more complicated than those for $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing. Reactions in Eqs. (8a) and (8b) also occur in reverse, and there are parallel processes involving $\nu_{e}$ and $\bar{\nu}_{e}$ as well:
$\nu_{e}+n$ $\rightleftharpoons p+e^{-},$ (9a)
$\bar{\nu}_{e}+p$ $\rightleftharpoons n+e^{+}.$ (9b)
Consequently, in order to treat the feedback on $V_{\nu}$ and $V_{\bar{\nu}}$ for $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing, we must include all the processes described above to follow the evolution of $Y_{n}$, $Y_{p}$, $Y_{\mu}$, $Y_{e}$, $Y_{\nu_{\mu}}$, and $Y_{\nu_{e}}$ simultaneously.
As mentioned in Section II, the particles of concern, $n$, $p$, $\mu^{\pm}$, $e^{\pm}$, $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$, are assumed to have Fermi-Dirac energy distributions characterized by the same temperature $T$ but species-specific chemical potentials. For fixed $T$ and $n_{b}$, the number fractions $Y_{n}$, $Y_{p}$, $Y_{\mu}$, $Y_{e}$, $Y_{\nu_{\mu}}$, and $Y_{\nu_{e}}$ are specified by the corresponding chemical potentials [see e.g., Eq. (2) for $Y_{\nu_{\mu}}$]. We use the following six constraints to determine these chemical potentials:
$Y_{n}+Y_{p}$ $=1,$ (10a)
$Y_{e}+Y_{\mu}$ $=Y_{p},$ (10b)
$\mu_{\nu_{e}}+\mu_{n}$ $=\mu_{p}+\mu_{e},$ (10c)
$Y_{p}$ $=\frac{\Gamma_{n\to p}}{\Gamma_{n\to p}+\Gamma_{p\to n}},$ (10d)
$\dot{Y}_{e}+\dot{Y}_{\nu_{e}}$ $=\Gamma_{L_{e}},$ (10e)
$\dot{Y}_{\mu}+\dot{Y}_{\nu_{\mu}}$ $=\Gamma_{L_{\mu}},$ (10f)
where $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ are the net rates for interconverting $n$ and $p$, and $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ are the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$ (see below). The constraints in Eqs. (10a), (10b), (10e), and (10f) correspond to conservation of baryon number, electric charge, and net lepton numbers of the electron and muon types. Because the reactions in Eqs. (9a) and (9b) are very efficient, we have assumed that chemical equilibrium is achieved among $n$, $p$, $e^{\pm}$, $\nu_{e}$, and $\bar{\nu}_{e}$ [Eq. (10c)], and that $Y_{p}$ takes the equilibrium value determined by $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ [Eq. (10d), see Appendix].
The net rates for interconverting $n$ and $p$ are
$\Gamma_{n\to p}$ $=\Gamma_{\nu_{e}n}+\Gamma_{e^{+}n}+\Gamma_{\nu_{\mu}n}+\Gamma_{\mu^{+}n},$ (11a)
$\Gamma_{p\to n}$ $=\Gamma_{\bar{\nu}_{e}p}+\Gamma_{e^{-}p}+\Gamma_{\bar{\nu}_{\mu}p}+\Gamma_{\mu^{-}p},$ (11b)
where for example, $\Gamma_{\nu_{e}n}$ is the rate per neutron for the forward reaction in Eq. (9a) (see Appendix for details). For the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$, we consider two scenarios to highlight the effects of sterile neutrino production and illustrate those of active neutrino diffusion. In the first scenario, we ignore the transport of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ by diffusion and take
$\Gamma_{L_{e}}$ $=0,$ (12a)
$\Gamma_{L_{\mu}}$ $=\dot{Y}_{\nu_{\mu}}^{\rm MSW}+\dot{Y}_{\nu_{\mu}}^{\rm coll}.$ (12b)
In the second scenario, we take
$\Gamma_{L_{e}}$ $=\dot{Y}_{\nu_{e}}^{\rm diff},$ (13a)
$\Gamma_{L_{\mu}}$ $=\dot{Y}_{\nu_{\mu}}^{\rm MSW}+\dot{Y}_{\nu_{\mu}}^{\rm coll}+\dot{Y}_{\nu_{\mu}}^{\rm diff},$ (13b)
where
$\dot{Y}_{\nu_{\beta}}^{\rm diff}=\frac{1}{6\pi G_{F}^{2}n_{b}r^{2}}\frac{\partial}{\partial r}\left(\frac{r^{2}}{n_{b}}\frac{\partial\mu_{\nu_{\beta}}}{\partial r}\right)\,,$ (14)
is the rate of change in $Y_{\nu_{\beta}}$ due to $\nu_{\beta}$ ($\beta=e,\ \mu$) diffusion (e.g., PhysRevD.108.063025 ).
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. The left (right) panels show the results without (with) diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Similar to Fig. 3, but for $m_{s}=10$ keV and $\sin^{2}2\theta=10^{-12}$.
IV Results
We now present example calculations of the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. We expect that in addition to the feedback effects discussed in Section III, such mixing would also affect the evolution of temperature $T$ and density $\rho$. As mentioned in the introduction, such dynamic effects can be properly quantified only by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in CCSN simulations. Our goal here, however, is to elucidate how such mixing enhances muonization. As a reasonable approximation, we limit the duration of our calculations so that the radial profiles of $T$ and $\rho$ can be considered as fixed in time. As the initial conditions, we take a snapshot at time post bounce $t_{\rm pb}=1$ s of a $20\,M_{\odot}$ CCSN model (with the SFHo nuclear equation of state and with muonization) provided by the Garching group 2017PhRvL.119x2702B ; Bollig2016 . Based on the comparison of the conditions at $t_{\rm pb}=1$ s and those at the next output instant $t_{\rm pb}=1.25$ s (see Fig. 2), we consider it reasonable to assume no evolution of $T(r)$ and $\rho(r)$ for a duration of $\Delta t=0.1$ s. We also limit our calculations to within the radius $r=14.7$ km, which corresponds to a temperature of $T=25.6$ MeV and a density of $\rho=8.01\times 10^{13}$ g cm-3, so that all active neutrinos are in the diffusion regime throughout the calculations.
Starting with the conditions at $t_{\rm pb}=1$ s and taking $T(r)$ and $\rho(r)$ as fixed in time, we follow the evolution of $Y_{n}$, $Y_{p}$, $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ by solving Eqs. (10a)–(10f) simultaneously. We carry out the calculations up to $t_{\rm pb}=1.1$ s in multiple time steps. For each step, we require that the results have numerically converged (with differences less than 1% when the calculations are repeated with half the step size). We have also checked that throughout the calculations, the changes in the total energy density and pressure of each radial zone are small (always less than 10% but typically few percent or less) as compared to the initial values of these quantities. For illustration, we show the results for $t_{\rm pb}=1.0025$, 1.01, 1.05, and 1.1 s.
We take $\delta m^{2}\approx m_{s}^{2}$ with $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$, which are suggested by interpreting the X-ray line emission near 3.55 keV from galaxy clusters as due to sterile neutrino decay Bulbul:2014sua ; Boyarsky:2014jta . For comparison, we use $m_{s}=10$ keV and $\sin^{2}2\theta=10^{-12}$, which are chosen for sterile neutrinos to make up all the dark matter Ng:2019gch ; Roach:2022lgo . For convenience, hereafter we refer to the two adopted sets of mixing parameters by the corresponding $m_{s}$ values. Using the rates $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ in Eqs. (12a) and (12b) [Eqs. (13a) and (13b)], we show the evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ in the left (right) panels of Fig. 3 for $m_{s}=7.1$ keV. Similar results for $m_{s}=10$ keV are shown in Fig. 4. The evolution of $Y_{n}$ and $Y_{p}$ (not shown) can be obtained from that of $Y_{e}$ and $Y_{\mu}$ through Eqs. (10a) and (10b).
IV.1 Evolution without diffusion of electron and muon neutrino lepton numbers
We first ignore diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ and discuss the results (left panels of Figs. 3 and 4) due to the production and escape of $\nu_{s}$ and $\bar{\nu}_{s}$ only [with $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ in Eqs. (12a) and (12b)]. For our adopted mixing parameters, the dominant production of sterile neutrinos is through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. We show the corresponding rate $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ for change in $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ in the top panels of Fig. 5. Initially $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ increases with radius as $\bar{\nu}_{e}$ of higher energy goes through MSW resonances at larger radii (bottom panels of Fig. 5). The resulting production and escape of $\bar{\nu}_{s}$ changes $Y_{\nu_{\mu}}$ directly and alter $Y_{\mu}$, $Y_{e}$, $Y_{\nu_{e}}$, $Y_{n}$, and $Y_{p}$ through those charged-current reactions intercoverting $n$ and $p$. Consequently, the potential $V_{\bar{\nu}}$ is changed, which in turn affects the subsequent MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. These feedback effects are clearly shown in Fig. 5. At the same radius, the resonant energy $E_{R}\approx\delta m^{2}/(2V_{\bar{\nu}})$ increases with time, mostly due to the increase of $Y_{\mu}$ and hence decrease of $V_{\bar{\nu}}\sim\sqrt{2}G_{F}n_{b}[(Y_{n}/2)-Y_{\mu}]$. The evolution of $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ is more complicated because it depends on both the resonant energy and the evolving energy distribution of $\bar{\nu}_{\mu}$ (due to the change of $\mu_{\nu_{\mu}}$). The general tendency is that $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ decreases with time, i.e., the feedback of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing tends to turn off such mixing (e.g., 2011PhRvD..83i3014R ; 2019JCAP…12..019S ; PhysRevD.108.063025 ). The cases of $m_{s}=7.1$ and 10 keV are similar except that the relevant resonant energy is higher (due to a higher $\delta m^{2}$) and the corresponding MSW conversion is less adiabatic (mostly due to a much smaller $\sin^{2}2\theta$) for the latter. The smaller $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ for $m_{s}=10$ keV leads to smaller effects on $Y_{\nu_{\mu}}$, $Y_{\mu}$, $Y_{e}$, and $Y_{\nu_{e}}$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: Evolution of the rate $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ and the resonant energy $E_{R}$ as functions of radius $r$ and time $t$ for the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ with $(m_{s}/{\rm keV},\sin^{2}2\theta)=(7.1,7\times 10^{-11})$ (left panels) and $(10,10^{-12})$ (right panels). Diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ is ignored.
As shown in the left panels of Figs. 3 and 4, there is little evolution of $Y_{e}$, the evolution of $Y_{\nu_{e}}$ is less than that of $Y_{\nu_{\mu}}$, and the evolution of $Y_{\mu}$ is the largest. The simultaneous changes of $Y_{\mu}$, $Y_{\nu_{\mu}}$, $Y_{\nu_{e}}$, and $Y_{e}$ tend to drive the corresponding particles towards chemical equilibrium. For illustration, we show in the left panel of Fig. 6 the evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ for the zone at $r=14.1$ km. Because $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})$ stays close to zero, chemical equilibrium among the corresponding particles always holds to good approximation. Note that although we assume $\mu_{\nu_{e}}+\mu_{n}=\mu_{p}+\mu_{e}$ in view of the fast rates of the reactions in Eqs. (9a) and (9b), we do not assume that the rates of the similar reactions involving $\mu^{\pm}$, $\nu_{\mu}$, and $\bar{\nu}_{\mu}$ are fast enough to ensure $\mu_{\nu_{\mu}}+\mu_{n}=\mu_{p}+\mu_{\mu}$. Instead, we assume that $Y_{p}$ takes the equilibrium value determined by the net rates of the reactions interconverting $n$ and $p$, which include the fast reactions in Eqs. (9a) and (9b). Our results show that the charged-current reactions intercoverting $n$ and $p$ by $\mu^{\pm}$, $\nu_{\mu}$, and $\bar{\nu}_{\mu}$ are sufficiently fast to establish $\mu_{\nu_{\mu}}+\mu_{n}\approx\mu_{p}+\mu_{\mu}$, which together with $\mu_{\nu_{e}}+\mu_{n}=\mu_{p}+\mu_{e}$, gives $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$.
Refer to captionRefer to caption
Figure 6: Evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ as functions of time $t$ for the zone at $r=14.1$ km. The left (right) panels show the results without (with) diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. Chemical equilibrium among the relevant particles is achieved to good approximation because $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$.
The relatively large change of $Y_{\mu}$ compared with much smaller changes of $Y_{\nu_{\mu}}$, $Y_{\nu_{e}}$, and $Y_{e}$ merits discussion. From $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$ and the chemical potentials shown in the left panel of Fig. 6, we expect similar changes in $\mu_{\mu}$ and $\mu_{\nu_{\mu}}$. However, for the conditions in the PNS, the increase of $Y_{\mu}$ with $\mu_{\mu}$ is much steeper than that of $Y_{\nu_{\mu}}$ with $\mu_{\nu_{\mu}}$ because of the muon rest mass (see Fig. 7). Therefore, the modest increase of $Y_{\nu_{\mu}}$ mainly due to the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ gives rise to the substantial increase of $Y_{\mu}$ through chemical equilibrium.
Refer to captionRefer to caption
Figure 7: Net number fractions $Y_{\mu}$ (left panel) and $Y_{\nu_{\mu}}$ (right panel) as functions of chemical potentials $\mu_{\mu}$ and $\mu_{\nu_{\mu}}$, respectively, for the zone at $r=14.1$ km with $T=28.9$ MeV and $\rho=9.99\times 10^{13}$ g cm-3. The solid points indicate the values at $t_{\rm pb}=1$ and 1.1 s from the calculations without diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. Because of the muon rest mass, the increase of $Y_{\mu}$ with $\mu_{\mu}$ is much steeper than that of $Y_{\nu_{\mu}}$ with $\mu_{\nu_{\mu}}$. Therefore, similar increases of $\mu_{\mu}$ and $\mu_{\nu_{\mu}}$ due to chemical equilibrium result in an increase of $Y_{\mu}$ much larger than that of $Y_{\nu_{\mu}}$.
IV.2 Evolution with diffusion of electron and muon neutrino lepton numbers
We find that diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ enhances the evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ (right panels of Figs. 3 and 4). The rate $\dot{Y}_{\nu_{e}}^{\rm diff}$ for change in $Y_{\nu_{e}}$ due to diffusion is $\lesssim 1$ s-1, and $\dot{Y}_{\nu_{\mu}}^{\rm diff}$ is $\lesssim 3$ s-1. These rates are comparable to $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ (see Fig. 5). In addition, $\dot{Y}_{\nu_{e}}^{\rm diff}$ and $\dot{Y}_{\nu_{\mu}}^{\rm diff}$ tend to increase with time, in contrast to the decreasing trend of $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$. Consequently, $Y_{\nu_{\mu}}$ reaches significantly higher values with diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$. Through chemical equilibrium, this increase of $Y_{\nu_{\mu}}$ drives the simultaneous increase of $Y_{\mu}$ similarly to the case without diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ (see Fig. 7), except that the change in $\mu_{\mu}$ is approximately the change in $\mu_{\nu_{\mu}}-\mu_{\nu_{e}}$ with significant evolution of $\mu_{\nu_{e}}$ (see the right panel of Fig. 6).
V Discussion and conclusions
We have presented the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS. In addition to the obvious feedback on the $\nu_{\mu}$ lepton number, we have included for the first time, the feedback on the composition of the PNS. For our adopted mixing parameters, which are consistent with current constraints, we find that sterile neutrinos are dominantly produced through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. This production and the subsequent escape of $\bar{\nu}_{s}$ increase the $\nu_{\mu}$ lepton number, which in turn enhances muonization mainly through $\nu_{\mu}+n\to p+\mu^{-}$ and changes the number fractions $Y_{e}$, $Y_{\nu_{e}}$, $Y_{p}$, and $Y_{n}$ through chemical equilibrium. We find that diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ enhances the evolution of the number fractions of all the particles. In particular, such diffusion further enhances muonization.
In our adopted CCSN model 2017PhRvL.119x2702B , $Y_{\mu}$ reaches the peak value of 0.0566 at $r=8.97$ km at $t_{\rm pb}=1$ s with standard physics. By including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\approx m_{s}^{2}\sim 10^{2}$ keV2, we have shown that enhanced muonization occurs in that model at larger radii beyond the original peak of $Y_{\mu}$ (Figs. 3 and 4). This enhancement is driven by the MSW resonances of $\bar{\nu}_{\mu}$ with $E_{R}\sim 4$–10 (8–20) MeV for $m_{s}=7.1$ (10) keV (Fig. 5), and is amplified by diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ (right panels of Figs. 3 and 4). While our treatment of such diffusion is approximate (e.g., with the same diffusion coefficient for all active neutrinos), it is adequate to demonstrate the above amplification effect and provide a reasonable estimate of this effect. In summary, we regard enhanced muonization as a robust result from $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS.
We have chosen to study the epoch of $t_{\rm pb}\approx 1$ s because the corresponding temperature and density profiles change so slowly that we can ignore such changes when studying the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing for a duration of 0.1 s. On the other hand, the PNS conditions do not differ greatly for $t_{\rm pb}\sim 0.4$–2 s and the conditions at $t_{\rm pb}\sim 0.1$ s are also not qualitatively different, which suggests that our results can be extended to earlier and later times. With the help of muonization, neutrino-driven explosion occurs at $t_{\rm pb}\sim 0.24$ s in our adopted CCSN model 2017PhRvL.119x2702B . We expect that when $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 are included in CCSN simulations, enhanced muonization similar to that found here would occur at the explosion epoch, potentially making the explosion easier. However, due to the intrinsically dynamic nature of the explosion, a definitive conclusion can only be reached by implementing such mixing in CCSN simulations.
Subsequent to the explosion, cooling of the PNS by active neutrino emission lasts up to $\sim 10$ s. Active neutrinos emitted from the PNS drive winds (e.g., 1996ApJ…471..331Q ) that may be an important source for heavy element nucleosynthesis (e.g., 1997ApJ…482..951H ). Further, this nucleosynthesis may be affected by collective oscillations of active neutrinos (see e.g., FISCHER2024104107 for a recent review) that are sensitive to the net lepton number carried by $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ fluxes PhysRevLett.125.251801 . With enhanced muonization due to $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the PNS, we expect that a significant net $\nu_{\mu}$ lepton number is emitted from the PNS, but note that the quantitative characteristics of active neutrino emission can only be obtained by including such mixing in CCSN simulations. We will explore the implications of enhanced muonization for the CCSN dynamics and collective neutrino oscillations in future works.
ACKNOWLEDGMENT
We sincerely thank Daniel Kresse and Thomas Janka for providing the CCSN model used in this work. This work is supported in part by the National Science Foundation (Grant No. PHY-2020275), the Heising-Simons Foundation (Grant 2017-228), and the U.S. Department of Energy (Grant No. DE-FG02-87ER40328). *
Appendix A Charged-current weak reaction rates
The evolution of $Y_{p}$ with time is governed by
$\dot{Y}_{p}$ $=\Gamma_{n\to p}Y_{n}-\Gamma_{p\to n}Y_{p}$
$=\Gamma_{n\to p}-(\Gamma_{n\to p}+\Gamma_{p\to n})Y_{p},$ (15)
where $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ are defined in terms of the rates for the charged-current weak reactions involving the electron and muon flavors [Eqs. (11a) and (11b)], and the second line follows from $Y_{n}+Y_{p}=1$. Regardless of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing, the rates for the charged-current weak reactions involving the electron flavor are very fast. Therefore, we can assume that $Y_{p}=\Gamma_{n\to p}/(\Gamma_{n\to p}+\Gamma_{p\to n})$ [Eq. (10d)] to very good approximation.
The rate for $\nu_{e}+n\to p+e^{-}$ is given by
$\Gamma_{\nu_{e}n}$ $=\frac{1}{n_{b}Y_{n}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{n}(E_{n})[1-f_{p}(E_{p})]$
×$\times\int_{0}^{\infty}\frac{E_{\nu_{e}}^{2}dE_{\nu_{e}}}{2\pi^{2}}f_{\nu_{e}}(E_{\nu_{e}})[1-f_{e^{-}}(E_{e^{-}})]\sigma_{\nu_{e}n}(E_{e^{-}}),$ (16)
where
$\sigma_{\nu_{e}n}(E_{e^{-}})=\frac{G^{2}_{F}\cos^{2}\theta_{c}}{\pi}E_{e^{-}}p_{e^{-}}$ (17)
is the cross section, $\theta_{c}$ denotes the Cabibbo angle with $\cos^{2}\theta_{c}=0.97$,
$f_{\alpha}(E_{\alpha})=\frac{1}{\exp[(E_{\alpha}-\mu_{\alpha})/T]+1}\,,$ (18)
is the occupation number of species $\alpha$ with $\alpha=n$, $p$, $\nu_{e}$, and $e^{-}$, and the chemical potential $\mu_{\alpha}$ includes the rest mass $m_{\alpha}$ of the species. In the above equations,
$E_{n,p}$ $=\sqrt{p^{2}+(m_{n,p}^{*})^{2}}-m_{n,p}^{*}+m_{n,p}+U_{n,p},$ (19)
$E_{e^{-}}$ $=\sqrt{p_{e^{-}}^{2}+m_{e}^{2}}$ (20)
$=E_{\nu_{e}}+m_{n}-m_{p}+U_{n}-U_{p},$ (21)
where for example, $m_{n}^{*}$ and $U_{n}$ are the effective mass and potential, respectively, for neutrons in the PNS (e.g., FISCHER2024104107 ).
The rate $\Gamma_{\nu_{\mu}n}$ for $\nu_{\mu}+n\to p+\mu^{-}$ can be obtained by replacing quantities for $\nu_{e}$ ($e^{-}$) with those for $\nu_{\mu}$ ($\mu^{-}$) in the equations for $\Gamma_{\nu_{e}n}$.
The rate for $\bar{\nu}_{e}+p\to n+e^{+}$ is given by
$\Gamma_{\bar{\nu}_{e}p}$ $=\frac{1}{n_{b}Y_{p}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{p}(E_{p})[1-f_{n}(E_{n})]$
×$\times\int_{E_{\bar{\nu}_{e}}^{\rm th}}^{\infty}\frac{E_{\bar{\nu}_{e}}^{2}dE_{\bar{\nu}_{e}}}{2\pi^{2}}f_{\bar{\nu}_{e}}(E_{\bar{\nu}_{e}})[1-f_{e^{+}}(E_{e^{+}})]\sigma_{\bar{\nu}_{e}p}(E_{e^{+}}),$ (22)
where $\sigma_{\bar{\nu}_{e}p}(E_{e^{+}})$ can be obtained by replacing $E_{e^{-}}$ ($p_{e^{-}}$) with $E_{e^{+}}$ ($p_{e^{+}}$) in Eq. (17) for $\sigma_{\nu_{e}n}(E_{e^{-}})$,
$E_{e^{+}}=E_{\bar{\nu}_{e}}+m_{p}-m_{n}+U_{p}-U_{n},$ (23)
and $E_{\bar{\nu}_{e}}^{\rm th}$ corresponds to $E_{e^{+}}=m_{e}$. Note that in the occupation number $f_{\bar{\alpha}}(E_{\bar{\alpha}})$ for an antiparticle $\bar{\alpha}$, the chemical potential is $\mu_{\bar{\alpha}}=-\mu_{\alpha}$.
The rate $\Gamma_{\bar{\nu}_{\mu}p}$ for $\bar{\nu}_{\mu}+p\to n+\mu^{+}$ can be obtained by replacing quantities for $\bar{\nu}_{e}$ ($e^{+}$) with those for $\bar{\nu}_{\mu}$ ($\mu^{+}$) in the equations for $\Gamma_{\bar{\nu}_{e}p}$.
The rate for $e^{+}+n\to p+\bar{\nu}_{e}$ is given by
$\Gamma_{e^{+}n}$ $=\frac{1}{n_{b}Y_{n}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{n}(E_{n})[1-f_{p}(E_{p})]$
×$\times\int_{0}^{\infty}\frac{p_{e^{+}}^{2}dp_{e^{+}}}{\pi^{2}}f_{e^{+}}(E_{e^{+}})[1-f_{\bar{\nu}_{e}}(E_{\bar{\nu}_{e}})]\sigma_{e^{+}n}(E_{\bar{\nu}_{e}}),$ (24)
where
$\sigma_{e^{+}n}(E_{\bar{\nu}_{e}})=\frac{G^{2}_{F}\cos^{2}\theta_{c}}{2\pi}E_{\bar{\nu}_{e}}^{2}$ (25)
is the cross section, and $E_{\bar{\nu}_{e}}=E_{e^{+}}+m_{n}-m_{p}+U_{n}-U_{p}$.
The rate $\Gamma_{\mu^{+}n}$ for $\mu^{+}+n\to p+\bar{\nu}_{\mu}$ can be obtained by replacing quantities for $e^{+}$ ($\bar{\nu}_{e}$) with those for $\mu^{+}$ ($\bar{\nu}_{\mu}$) in the equations for $\Gamma_{e^{+}n}$.
The rate for $e^{-}+p\to n+\nu_{e}$ is given by
$\Gamma_{e^{-}p}$ $=\frac{1}{n_{b}Y_{p}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{p}(E_{p})[1-f_{n}(E_{n})]$
×$\times\int_{E_{e^{-}}^{\rm th}}^{\infty}\frac{p_{e^{-}}^{2}dp_{e^{-}}}{\pi^{2}}f_{e^{-}}(E_{e^{-}})[1-f_{\nu_{e}}(E_{\nu_{e}})]\sigma_{e^{-}p}(E_{\nu_{e}}),$ (26)
where $\sigma_{e^{-}p}(E_{\nu_{e}})$ can be obtained by replacing $E_{\bar{\nu}_{e}}$ with $E_{\nu_{e}}$ in Eq. (25) for $\sigma_{e^{+}n}(E_{\bar{\nu}_{e}})$,
$E_{\nu_{e}}=E_{e^{-}}+m_{p}-m_{n}+U_{p}-U_{n},$ (27)
and $E_{e^{-}}^{\rm th}$ corresponds to $E_{\nu_{e}}=0$.
The rate $\Gamma_{\mu^{-}p}$ for $\mu^{-}+p\to n+\nu_{\mu}$ can be obtained by replacing quantities for $e^{-}$ ($\nu_{e}$) with those for $\mu^{-}$ ($\nu_{\mu}$) in the equations for $\Gamma_{e^{-}p}$.